# Infinite magneto-resistance and bipolar effect in spin valves driven by spin batteries

### v1

@@ 1 Introduction
The IMR does not rely on specific materials such as half-metals to increase GMR (Heusler alloys as in [5]) or coherent tunneling through MgO barriers [6, 7]. It consists in achieving a zero signal for a system with 2 magnetically active parts (like two ferromagnetic layers in a spin valve) so that the optimistic (resp. pessimistic) MR ratio tends to infinity (resp. unity), hence the naming of the effect as IMR. Its main characteristics are: (1) the use of a spin battery instead of a charge battery; (2) the reliance on symmetry to get rid of the offsets preventing a bipolar response in Johnson BST; (3) the reliance on Johnson-Silsbee charge-spin coupling [8] to measure the effect. Of course, the effect is curtailed by asymmetries which always exist so that the MR ratio can never be infinite and it will be one of the objectives of the paper to quantify to impact of slight asymmetries and show that the MR ratio can still be very large.
@@
What is of paramount importance is to have symmetrical setups which will allow to get rid by construction of any offsets. A simple recipe is to have two copies of the same system as in Fig. 1-a: the offsets obviously cancel out. This is an example of what we will christen as ’twin systems’.
Figure 1: (a) The voltage measured across two copies of the same system (and batteries) vanishes by symmetry. (b) FNF valve connected to a spin battery and a voltmeter. If the combined system is symmetrical under left-right exchange, then the voltage must vanish by symmetry. The layers can be viewed as parallel to the figure plane (as in non-local geometries) or to be perpendicular. SB: spin battery. (c) Same as (b) but depicted with an FNF nanopillar playing the role of the SB. (d) CPP trilayer driven symmetrically by two spin batteries. For parallel magnetizations, the voltage again vanishes (for symmetric probes). The black arrows depict the magnetization directions.
Another example of a ’twin system’ is given in Fig. 1-b. It is an $FNF$ valve in a non-local geometry with identical ferromagnets and layers parallel to the figure plane; it could also be considered to be a trilayer with layers perpendicular to the figure plane. A spin battery injects a spin current through the intermediate $N$ layer. The charge voltage across the trilayer obviously vanishes when the ferromagnets have parallel magnetizations but is non-zero for anti-parallel ones. In practice the vanishing of the signal is constrained by the degree of symmetry. Suppose for instance that the widths of the ferromagnets are equal to $l_{m}\pm\delta l$. Then we expect a scaling:
@@
There are other possible setups than ’twin setups’ which will exhibit IMR and bipolarity, as a result of symmetry. For instance consider a CPP symmetric trilayer where instead of a charge battery, one connects the layers to two spin batteries as in Fig. 1-d. (The CPP qualifier, current perpendicular to plane, comes from the fact that if one replaced the voltage probes by current ones, the charge current would be perpendicular to the layers.) This setup exhibits IMR for identical incoming spin currents and parallel magnetizations. We will discuss later a similar setup with opposite spin currents; one then has IMR for antiparallel magnetizations.
@@
The latter case is actually easy to understand in terms of a resistor analogy. We consider a spin accumulation source for the spin battery (a spin current source will yield similar results); in this example we will measure a charge current. In the resistor analogy for a CPP symmetric spin valve with anti-parallel magnetizations driven by a charge battery, currents flowing in each spin channel are equal $I_{\uparrow}=I_{\downarrow}$. The key observation is to ask what would happen if the spin channels had a reverse bias (opposite voltages) as in Fig. 2-(b) : the currents in each spin channel are now opposite $I_{\uparrow}=-I_{\downarrow}$; obviously no charge current can thus flow out of the device (since $I_{c}=I_{\uparrow}+I_{\downarrow}$) in contrast to Fig.2-(a) (parallel case). But a reverse bias is precisely what a spin accumulation source can achieve: indeed a spin battery splits the chemical potentials for each spin channel $V_{\uparrow}\neq V_{\downarrow}$ and additionally $V_{\uparrow}=-V_{\downarrow}$ since $V_{c}=0$ in the absence of a charge voltage applied to the system. Replacing a charge battery with a spin battery therefore generates reversed biases and opposite currents in each spin channel (for a symmetric spin valve), so that in the anti-parallel state, there is no outgoing charge current: the corresponding ratio $\left(I_{c}^{AP}-I_{c}^{P}\right)/I_{c}^{AP}$ goes to infinity (IMR effect).
Figure 2: In the resistor network model, symmetric CPP spin valve driven by a spin battery which injects a spin current and induces spin accumulation. The latter is equivalent to reverse biasing the two spin channels. The currents (in electrical units) in each spin channel are $I_{\uparrow}$ and $I_{\downarrow}$. (a) Parallel magnetizations for ferromagnets $F1$ and $F2$ : non-vanishing charge current $I_{c}^{P}\neq 0$. (b) Anti-parallel magnetizations: current $I_{c}$ at detector is vanishing since the currents in each channel are opposite. Therefore one has an infinite resistance state.
2 Response function, BE & IMR.
2.1 Trans-resistance.
@@
$r_{cs}=\left(\frac{\partial V_{c}}{\partial I_{s}}\right)_{I_{c}=0}.$ (5)
It has the dimension of a resistance. The basic geometries we consider are therefore 3 terminal ones, two terminals used for the voltage measurement and one terminal connected to the spin battery. We will make the hypothesis that the spin battery is a perfect spin current source which outputs a constant spin current $I_{s}$, whatever the load it is attached to. This is not a very stringent condition. It is easy to take into account spin backflow into the battery through the internal spin resistance concept [26], which in practice will translate into a reduced response. Let us suppose that the load comprises $N=2$ magnetically active regions, whose magnetizations are collinear (but maybe have different directions) and can be controlled independently. One can write therefore: $r_{cs}=r_{cs}\left(\sigma_{1},\;\sigma_{2}\right)$ where $\left\{\sigma_{1}=\pm 1,\;\sigma_{2}=\pm 1\right\}$ index the magnetizations of the $2$ regions relative to a given quantization axis. $r_{cs}$ is simply a multilinear polynomial in $\sigma_{1}$ and $\sigma_{2}$:
@@
Since the condition for IMR also allows the BE, let us compare the two effects for the setups of Fig. 1-a, 1-b. IMR is observed by switching a single magnetization while bipolarity requires switching two magnetizations, which may or may not be possible or desireable in a given setup. The amplitude of the IMR effect is $\left|r_{cs}^{AP}-r_{cs}^{P}\right|=2r_{1}$, while the bipolar effect has an amplitude which is twice as large. There is therefore a trade-off between the two effects: IMR is smaller but easier to observe, requiring a single magnetization to be switched, while bipolarity has the larger signal but requires both magnetizations to be switched.
@@
3 Example setups.
3.1 Twin setup with shared spin battery.
@@
$r_{cs}^{I}$ $=\left(P_{F1}-P_{F2}\right)\>\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right]\>\frac{r_{F}}{2}+\delta l\>\left(P_{F1}+P_{F2}\right)\>\frac{r_{F}}{2}\>\alpha,$ (11)
@@
where $\alpha$ and $\beta$ are factors which depend on the geometry and the various spin resistances (for their exact expressions, we refer to [27]).
$r_{cs}^{III}$ $=\left(P_{c1}^{\prime}-P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)+\delta l\>\left(P_{c1}^{\prime}+P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$ (13)
@@
It is also important to have an idea of the range of the HIGH signal. Let’s use MSU results for CoCu at $4.2\>K$ and for simplicity assume that $P_{c}\sim P_{c}^{\prime}=0.77\pm 0.05$, $P_{F}=0.46\pm 0.05,$ $2r_{c}A\sim 2r_{c}^{\prime}A=1.02\pm 0.04\>f\Omega m^{2}$, $t_{Co}=2-10\,nm,\>l_{Co}>60\>nm$ and $\rho_{F}(Co)=75\pm 5\>n\Omega m,$ $\rho_{N}(Cu)=6\pm 1\>n\Omega m.$ Then $X\sim 1$, $f_{m}\sim 1$. So for AP ferromagnets:
@@
3.2 Twins with individual spin batteries.
@@
3.3 Symmetric CPP spin valves operated by two spin batteries. Parallel IMR.
@@
It is therefore advantageous to engineer symmetrical systems which can have reproducible behaviours. Consider for instance Fig.1-d which shows a spin valve in a CPP geometry (if the voltmeter was shorted, a charge current would flow in a direction orthogonal to the layers). Note that the setup in Fig. 1-d could represent also a nano-pillar or a non-local geometry in addition to the usual multi-layer thin films. The spin current from the batteries are oriented to be incoming in the load (outgoing from the batteries). If the ferromagnetic layers and the spin batteries are identical, there is a plane of symmetry $\mathcal{P}$ in the paramagnetic layer parallel to the layers. Obviously $V_{c}=0$ when the magnetizations are parallel. One has a parallel IMR. If one starts with antiparallel magnetizations and switches both of them, one gets a bipolar effect.
@@
and a bipolar effect for antiparallel magnetizations when flipping both of them for identical currents: $V_{c}\left(\sigma_{1},\>-\sigma_{1}\right)=-V_{c}\left(-\sigma_{1},\>\sigma_{1}\right).$ Detailed calculations on an FNF trilayer exhibiting such a parallel IMR are given elsewhere [27]. In the limit of infinite ferromagnets, one finds
@@
3.4 Anti-parallel IMR in CPP MTJ & FNF trilayer.
It is also possible to engineer an anti-parallel IMR. Fig.3-a shows such a setup. A CPP trilayer is connected symmetrically to two spin batteries which inject opposite spin currents. If the voltage probes are also symmetrical, the charge voltage must vanish as shown in Fig.3. Indeed let us rotate the setup by an angle $\pi$ around an axis $\mathcal{P}$ in the middle of the layers (Fig.3-d): this exchanges the spin batteries and the ferromagnets and the voltage drop gets reversed $V_{c}\longrightarrow-V_{c}$. But one could also rotate the setup by an angle $\pi$ around an axis $\mathcal{F}$ splitting the layers symmetrically (Fig.3-b): all the magnetizations get reversed; the spin current reference axis is also reversed but the charge voltage is obviously unchanged. But a spin current $+I_{s}$ against a given magnetization reference axis $\mathbf{M}$ is equivalent to a spin current $-I_{s}$ with a reversed reference axis $-\mathbf{M}$. One therefore ends up with Fig.3-c which is equivalent to Fig.3-d apart from an opposite voltage. This implies that $V_{c}=0$. The obvious issue with such a setup is that both the spin batteries and the voltage probes must be located at the same positions. This is however easy to relax by doubling the voltage probes and positioning them in a symmetric manner with respect to symmetry plane $\mathcal{F}$ (see Fig.3-e); one would then measure two voltages drops $V_{c1}$ and $V_{c2}$ which might not be offset-free but their average will be: $V_{c}=\left(V_{c1}+V_{c2}\right)/2$.
Figure 3: (a) CPP symmetric spin valve connected to two spin batteries injecting opposite spin currents $I_{s}$ into the ferromagnets. The black arrows show the magnetization direction in each layer ($F1$or $F2$) or the reference direction used to measure the spin current (in the spin batteries). (b) Setup rotated around a symmetry axis $\mathcal{D}$ assumed to exist. (b) Previous setup where the spin currents take opposite values if the reference axis is flipped in the spin batteries. (d) Setup rotated by an angle $\pi$ around a symmetry axis $\mathcal{P}$ assumed to exist in the middle. Comparing (c) and (d) shows they’re identical setups provided the ferromagnets $F1$ $F2$ are identical (for symmetrical voltage probes), so that for anti-parallel magnetization, there is no voltage drop across terminals $T_{1}$ and $T_{2}$. (e) Instead of two voltage probes in the middle of the layers, it might be easier to use 2 sets of voltage probes positioned symmetrically with respect to axis $\mathcal{F}$.
@@
Let us consider layers which are not infinitely thin; for instance, we assume equal widths $d_{1}=d_{2}=5\>nm$ up to a monolayer; this means $l_{1/2}=d_{1/2}/l_{F}=0.125$. Using the same figures as above numbers, a more realistic calculation [27] shows that the spin current asymmetry can usually be neglected and leads to $TMR\sim 10^{5}\,\%$, which is still extremely large.
For an FNF trilayer in the same geometry, in the limit of thin layers, one gets a similar MR ratio. The ratio is however much higher for an MTJ because $r_{c}$ is considerably larger by a factor $10^{3}-10^{5}$. For all metallic spin valves, a spin battery based on ferromagnetic resonance would be able to achieve HIGH signals in the $1-100\>\mu V$ range. By using an MTJ, since the resistance is multiplied by a factor $10^{3}-10^{5}$, this implies that $mV$ signals are very achievable which would make such a spin battery driven MTJ quite useable.
@@
4 Conclusion.
We have introduced several spin valve setups which due to symmetry can exhibit both the bipolar effect of Johnson spin transistor and an IMR. An obvious application of IMR and BE would be more sensitive magnetic sensors so an experimental confirmation of these predictions would be very useful. An essential component for the occurence of both BE and IMR is to replace charge batteries by spin batteries since this allows the occurence in response functions of odd terms $\propto sgn\left(M_{z}\right)$. Since real systems are never perfectly symmetric, we have also quantified the impact of length & spin current asymmetries. The MR ratio can still reach high ranges $10-10^{3}$. The simplest system to observe the BE is the twin system with shared spin battery; for observation of IMR, it is better to test on a twin system with independent spin batteries. The CPP MTJ with two spin batteries is quite promising in comparison with the CPP FNF trilayer due to the much larger HIGH signal which should make the $mV$ range quite reachable. Extensions of this work to non-collinear systems with symmetries, to CIP setups, to parallel IMR in CPP systems and inclusion of non-ideal spin batteries are discussed elsewhere [27].

### v3

@@ I Introduction
The IMR does not rely on specific materials such as half-metals to increase GMR (Heusler alloys as in Kimura et al[5]) or coherent tunneling through MgO barriers [6, 7]. It consists in achieving a zero signal for a system with 2 magnetically active parts (like two ferromagnetic layers in a spin valve) so that the optimistic (resp. pessimistic) MR ratio tends to infinity (resp. unity), hence the naming of the effect as IMR. Its main characteristics are: (1) the use of a spin battery instead of a charge battery; (2) the reliance on symmetry to get rid of the offsets preventing a bipolar response in Johnson BST; (3) the reliance on Johnson-Silsbee charge-spin coupling [8] to measure the effect. Of course, the effect is curtailed by asymmetries which always exist so that the MR ratio can never be infinite and it will be one of the objectives of the paper to quantify to impact of slight asymmetries and show that the MR ratio can still be very large.
@@
What is of paramount importance is to have symmetrical setups which will allow to get rid by construction of any offsets. A simple recipe is to have two copies of the same system as in Fig. 1-a: the offsets obviously cancel out. This is an example of what we will christen as ’twin systems’. In Fig. 1, the ferromagnets can be thought of being in a non-local geometry. But they could also be thin films as in Fig. 2 -a
Figure 1: (a) The voltage measured across two copies of the same system (and batteries) vanishes by symmetry. (b) If one ferromagnet is flipped, one gets a non-vanishing voltage. (c) If both ferromagnets are flipped from (a), the voltage still vanishes. (d) The voltage is now the opposite of (b) since (d) follows from (b) by flipping both magnetizations, which amount to exchanging both systems.
Another example of a ’twin system’ is given in Fig. 2-b. It is an $FNF$ valve in a non-local geometry with identical ferromagnets and layers parallel to the figure plane; it could also be considered to be a trilayer with layers perpendicular to the figure plane. A spin battery injects a spin current through the intermediate $N$ layer. The charge voltage across the trilayer obviously vanishes when the ferromagnets have parallel magnetizations but is non-zero for anti-parallel ones.
Figure 2: (Color online) (a) The voltage measured across two copies of the same system (and batteries) vanishes by symmetry. (b) FNF valve connected to a spin battery and a voltmeter. If the combined system is symmetrical under left-right exchange, then the voltage must vanish by symmetry. The layers can be viewed as parallel to the figure plane (as in non-local geometries) or to be perpendicular. SB: spin battery. (c) Same as (b) but depicted with an FNF nanopillar playing the role of the SB. (d) CPP trilayer driven symmetrically by two spin batteries. For parallel magnetizations, the voltage again vanishes (for symmetric probes). (e) Twin system with an angular offset. The black arrows depict the magnetization directions.
In practice the vanishing of the signal is constrained by the degree of symmetry. Suppose for instance that the widths of the ferromagnets are equal to $l_{m}\pm\delta l$. Then we expect a scaling:
@@
There are other possible setups than ’twin setups’ which will exhibit IMR and bipolarity, as a result of symmetry. For instance consider a CPP symmetric trilayer where instead of a charge battery, one connects the layers to two spin batteries as in Fig. 2-d. (By CPP, current perpendicular to plane, we mean that if one replaced the voltage probes by current probes, the charge current would flow perpendicular to the layers.) This setup exhibits IMR for identical incoming spin currents and parallel magnetizations (parallel IMR). We will discuss later a similar setup with opposite spin currents; one then has IMR for antiparallel magnetizations (AP IMR). Note also that, as shown on 2-e, the two copies of the same system can have arbitrary relative orientations provided the voltage measurement probes remain symmetric.
Figure 3: (Color online) Resistor model of a spin valve driven by a spin current source. The charge voltage is $V_{c}=\left(V_{\uparrow}+V_{\downarrow}\right)/2$. (a) $V_{c}\neq 0$ when ferromagnets $F1$ and $F2$ have parallel magnetizations. (b) But for anti-parallel magnetizations, the voltage $V_{c}$ obviously vanishes since $V_{\uparrow}=\left(r+R\right)\>I_{s}/2=-V_{\downarrow}$.
@@
The AP IMR for CPP spin valve is actually easy to understand in terms of a resistor analogy with a single spin battery (due to current conservation on each spin channel, a second spin battery is irrelevant, see Fig.4-a). We consider a spin current source in a two-channel model ( Fig. 3) : this is tantamount to having opposite charge current sources for each spin channel. In the resistor analogy for a CPP symmetric spin valve with anti-parallel magnetizations driven by a charge battery, the currents flowing in each spin channel would be equal $I_{\uparrow}=I_{\downarrow}$ and the voltages for each spin channel would be identical in the anti-parallel case $V_{\uparrow}=V_{\downarrow}$ In the antiparallel case for a spin current source driving the resistors ( Fig.3-(b) ), since the currents are opposite $I_{\uparrow}=-I_{\downarrow}=I_{s}/2$, the channel voltages are also opposite $V_{\uparrow}=-V_{\downarrow}$ so that $V_{c}=0$. The charge voltage divided by the spin current has the dimension of a resistance $R=V_{c}/I_{s}$. The corresponding MR ratio $\left(R^{AP}-R^{P}\right)/R^{AP}$ goes to infinity (IMR effect). A similar resistor analogy can be drawn for parallel IMR for CPP spin valves but requires to take into account spin flips and transfer from one spin channel to the other. See Fig.4-b.
Figure 4: (Color online) (a) Resistor analogy for AP IMR of CPP spin valve connected to two spin batteries. Since spin flip is neglected, the spin current source on the right can be removed because it is redundant. (b) The resistor analogy for parallel IMR of CPP spin valve connected to 2 spin current sources requires addition of spin mixing either in bulk and/or at the interface between the ferromagnets. In the resistor circuit, the resistance $r^{\prime}$ ensures the current to be in a closed loop while no current flows through resistances $r$ and $R$. Therefore the charge voltage $V_{c}=\left(V_{\uparrow}+V_{\downarrow}\right)/2$ vanishes. The same conclusion is reached by adding a mixing resistance at the interface between the ferromagnets.
II Response function, BE & IMR.
II.1 Trans-resistance.
@@
$r_{cs}=\left(\frac{\partial V_{c}}{\partial I_{s}}\right)_{I_{c}=0},$ (5)
where the spin current is defined in electrical units as $I_{s}=I_{\uparrow}-I_{\downarrow}$ for collinear setups. It has the dimension of a resistance. The basic geometries we consider are therefore 3 terminal ones, two terminals used for the voltage measurement and one terminal connected to the spin battery. We will make the hypothesis that the spin battery is a perfect spin current source which outputs a constant spin current $I_{s}$, whatever the load it is attached to. This is not a very stringent condition. It is easy to take into account spin backflow into the battery through the internal spin resistance concept [26], which in practice will translate into a reduced response. Let us suppose that the load comprises $N=2$ magnetically active regions, whose magnetizations are collinear (but maybe have different directions) and can be controlled independently. One can write therefore: $r_{cs}=r_{cs}\left(\sigma_{1},\;\sigma_{2}\right)$ where $\left\{\sigma_{1}=\pm 1,\;\sigma_{2}=\pm 1\right\}$ index the magnetizations of the $2$ regions relative to a given quantization axis. $r_{cs}$ is simply a multilinear polynomial in $\sigma_{1}$ and $\sigma_{2}$:
@@
Since the condition for IMR also allows the BE, let us compare the two effects for the setups of Fig. 2-a, 2-b. IMR is observed by switching a single magnetization while bipolarity requires switching two magnetizations, which may or may not be possible or desireable in a given setup. The amplitude of the IMR effect is $\left|r_{cs}^{AP}-r_{cs}^{P}\right|=2r_{1}$, while the bipolar effect has an amplitude which is twice as large. There is therefore a trade-off between the two effects: IMR is smaller but easier to observe, requiring a single magnetization to be switched, while bipolarity has the larger signal but requires both magnetizations to be switched.
@@
III Example setups.
III.1 Twin setup with shared spin battery.
@@
$r_{cs}^{I}$ $=\left(P_{F1}-P_{F2}\right)\>\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right]\>\frac{r_{F}}{2}$
$+\delta l\>\left(P_{F1}+P_{F2}\right)\>\frac{r_{F}}{2}\>\alpha,$ (11)
@@
where $\alpha$ and $\beta$ are factors which depend on the geometry and the various spin resistances, namely
$\alpha$ $=\left[\cosh_{m}-\left(f_{m}^{2}-1\right)\left(\cosh_{m}-1\right)-f_{m}\sinh_{m}\right]$
$+\frac{f_{m}^{2}-1}{\frac{r_{c}}{r_{F}}+f_{m}}\left[-f_{m}\left(\cosh_{m}-1\right)+\sinh_{m}\right],$ (13)
$\beta$ $=\frac{f_{m}^{2}-1}{\frac{r_{c}}{r_{F}}+f_{m}}.$ (14)
$r_{cs}^{III}$ $=\left(P_{c1}^{\prime}-P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$ (15)
$+\delta l\>\left(P_{c1}^{\prime}+P_{c2}^{\prime}\right)\>\frac{r_{c}^{\prime}}{2}\>\left(\cosh_{m}-f_{m}\sinh_{m}\right)$
@@
It is also important to have an idea of the range of the HIGH signal (the LOW signal being close to 0). Let’s use MSU results for CoCu at $4.2\>K$ and for simplicity assume that $P_{c}\sim P_{c}^{\prime}=0.77\pm 0.05$, $P_{F}=0.46\pm 0.05,$ $2r_{c}A\sim 2r_{c}^{\prime}A=1.02\pm 0.04\>f\Omega m^{2}$, $t_{Co}=2-10\,nm,\>l_{Co}>60\>nm$ and $\rho_{F}(Co)=75\pm 5\>n\Omega m,$ $\rho_{N}(Cu)=6\pm 1\>n\Omega m.$ Then $X\sim 1$, $f_{m}\sim 1$. So for AP ferromagnets:
@@
III.2 Twins with individual spin batteries.
@@
III.3 Symmetric CPP spin valves operated by two spin batteries. Parallel IMR.
@@
It is therefore advantageous to engineer symmetrical systems which can have reproducible behaviours. Consider for instance Fig.2-d which shows a spin valve in a CPP geometry (if the voltmeter was shorted, a charge current would flow in a direction orthogonal to the layers). Note that the setup in Fig. 2-d could represent also a nano-pillar or a non-local geometry in addition to the usual multi-layer thin films. The spin current from the batteries are oriented to be incoming in the load (outgoing from the batteries). If the ferromagnetic layers and the spin batteries are identical, there is a plane of symmetry $\mathcal{P}$ in the paramagnetic layer parallel to the layers. Obviously $V_{c}=0$ when the magnetizations are parallel. One has a parallel IMR. If one starts with antiparallel magnetizations and switches both of them, one gets a bipolar effect. See Fig.4 for the resistor analogy.
@@
and a bipolar effect for antiparallel magnetizations when flipping both of them for identical currents: $V_{c}\left(\sigma_{1},\>-\sigma_{1}\right)=-V_{c}\left(-\sigma_{1},\>\sigma_{1}\right).$ Detailed calculations on an FNF trilayer exhibiting such a parallel IMR will be given elsewhere. In the limit of infinite ferromagnets, one finds in terms of a mean spin current $I_{s,0}^{m}=\left(I_{s,L}+I_{s,R}\right)/2$ and a spin current asymmetry $\delta I_{s}=\left(I_{s,L}-I_{s,R}\right)/2$:
@@
(Length asymmetry drops out for infinite systems.)
III.4 Anti-parallel IMR in CPP MTJ & FNF trilayer.
It is also possible to engineer an anti-parallel IMR (see Fig.3 for the resistor analogy explaining it). Fig.5-a shows such a setup. A CPP trilayer is connected symmetrically to two spin batteries which inject opposite spin currents. If the voltage probes are also symmetrical, the charge voltage must vanish as shown in Fig.5. Indeed let us rotate the setup by an angle $\pi$ around an axis $\mathcal{P}$ in the middle of the layers (Fig.5-d): this exchanges the spin batteries and the ferromagnets and the voltage drop gets reversed $V_{c}\longrightarrow-V_{c}$. But one could also rotate the setup by an angle $\pi$ around an axis $\mathcal{F}$ splitting the layers symmetrically (Fig.5-b): all the magnetizations get reversed; the spin current reference axis is also reversed but the charge voltage is obviously unchanged. But a spin current $+I_{s}$ against a given magnetization reference axis $\mathbf{M}$ is equivalent to a spin current $-I_{s}$ with a reversed reference axis $-\mathbf{M}$. One therefore ends up with Fig.5-c which is equivalent to Fig.5-d apart from an opposite voltage. This implies that $V_{c}=0$. The obvious issue with such a setup is that both the spin batteries and the voltage probes must be located at the same positions. This is however easy to relax by doubling the voltage probes and positioning them in a symmetric manner with respect to symmetry plane $\mathcal{F}$ (see Fig.5-e); one would then measure two voltages drops $V_{c1}$ and $V_{c2}$ which might not be offset-free but their average will be: $V_{c}=\left(V_{c1}+V_{c2}\right)/2$.
Figure 5: (Color online) (a) CPP symmetric spin valve connected to two spin batteries injecting opposite spin currents $I_{s}$ into the ferromagnets. The black arrows show the magnetization direction in each layer ($F1$or $F2$) or the reference direction used to measure the spin current (in the spin batteries). (b) Setup rotated around a symmetry axis $\mathcal{D}$ assumed to exist. (b) Previous setup where the spin currents take opposite values if the reference axis is flipped in the spin batteries. (c) Setup rotated by an angle $\pi$ around a symmetry axis $\mathcal{P}$ assumed to exist in the middle. (d) is (b) with reference directions flipped in the spin batteries (which changes the sign of spin currents). Comparing (c) and (d) shows they’re identical setups provided the ferromagnets $F1$ $F2$ are identical (for symmetrical voltage probes), so that for anti-parallel magnetization, there is no voltage drop across terminals $T_{1}$ and $T_{2}$. (e) Instead of two voltage probes in the middle of the layers, it might be easier to use 2 sets of voltage probes positioned symmetrically with respect to axis $\mathcal{F}$. The previous discussion then applies to $V_{c}=\left(V_{c1}+V_{c2}\right)/2$.
@@
Let us consider layers which are not infinitely thin; for instance, we assume equal widths $d_{1}=d_{2}=5\>nm$ up to a monolayer; this means $l_{1/2}=d_{1/2}/l_{F}=0.125$. Using the same figures as above numbers, a more realistic calculation (which will be detailed elsewhere) shows that the spin current asymmetry can usually be neglected and leads to $TMR\sim 10^{5}\,\%$, which is still extremely large.
For an FNF trilayer in the same geometry, in the limit of thin layers, one gets a similar MR ratio. The ratio is however much higher for an MTJ because $r_{c}$ is considerably larger, by a factor $10^{3}-10^{5}$. For all metallic spin valves, a spin battery based on ferromagnetic resonance would be able to achieve signals in the $1-100\>\mu V$ range. By using an MTJ, since the resistance is multiplied by a factor $10^{3}-10^{5}$, this implies that $mV$ signals are very achievable which would make such a spin battery driven MTJ quite useable.
@@
IV Conclusion.
We have introduced several spin valve setups which due to symmetry can exhibit both the bipolar effect of Johnson spin transistor and an IMR. An obvious application of IMR and BE would be more sensitive magnetic sensors so an experimental confirmation of these predictions would be very useful. An essential component for the occurence of both BE and IMR is to replace charge batteries by spin batteries since this allows the occurence in response functions of odd terms $\propto sgn\left(M_{z}\right)$. Since real systems are never perfectly symmetric, we have also quantified the impact of length & spin current asymmetries. The MR ratio can still reach high ranges $10-10^{3}$. The simplest system to observe the BE is the twin system with shared spin battery; for observation of IMR, it is better to test on a twin system with independent spin batteries. The CPP MTJ with two spin batteries is quite promising in comparison with the CPP FNF trilayer due to the much larger HIGH signal which should make the $mV$ range quite reachable. Extensions of this work to non-collinear systems with symmetries, to CIP setups, to parallel IMR in CPP systems and inclusion of non-ideal spin batteries are discussed elsewhere.